what is the vertex of the parabola y=x^2-36x+234
Answer by Alan3354(69443)

what is the vertex of the parabola y=x^2-36x+234
------
It's on the LOS, Line of Symmetry x = -b/2a
x = 36/2 = 18
----
y = 18^2 - 36*18 + 234 = -90
Vertex at (18,-90)